Show that the functions f: R -->1 , infinity -->R , defined by : x=3^2x +1 , x= log x-1 are inverse of one another.
1
Expert's answer
2021-09-17T03:50:21-0400
"f(x)=3^{2x}+1"

"Domain: (-\infin, \infin)"

"Range: (1, \infin)"

Replace "f(x)" by "y"


"y=3^{2x}+1"

Interchange "x" and "y"


"x=3^{2y}+1"

Solve for "y"


"3^{2y}=x-1"

"y=\dfrac{1}{2}\log_3(x-1)"

Replace "y" by "f^{-1}(x)"


"f^{-1}(x)=\dfrac{1}{2}\log_3(x-1)"

"Domain: (1, \infin)"

"Range: (-\infin, \infin)"

The functions "f(x)=3^{2x}+1" and "f^{-1}(x)=\dfrac{1}{2}\log_3(x-1)" are inverse of one another.